OpenStudy (macabbey101):

HELP PLEASE I WILL MEDAL! 2|x+7|+9=4|x+7|+3

OpenStudy (anonymous):

2x+14+9=x+3

OpenStudy (anonymous):

combine your x's and like terms

OpenStudy (anonymous):

2x=x-20

OpenStudy (anonymous):

subtract x form both sides for x=-20

OpenStudy (anonymous):

make sense?